Three resistors R1= 3.5 ohm, R2=2* 3.5 ohm and R3= 4.4 ohm are connected in parallel. What is the equivalent resistance?
Answer #1

let their equivalent resistance be R'.

\frac{1}{R'}=\frac{1}{R1}+\frac{1}{R2}+\frac{1}{R3}

\Rightarrow R' =\frac{R1\times R2\times R3}{R1+R2+R3}

So R' = 7.235ohm